OpenStudy (anonymous):

f(x) = 4x - 3 and g(x) = solve for g(f(2))

OpenStudy (anonymous):

@Cosmichaotic can u plz help!!!

OpenStudy (anonymous):

ok,g(x) is equal to what?

OpenStudy (anonymous):

g(x) = 2x-1/3

OpenStudy (anonymous):

ok,first find f(2)

OpenStudy (anonymous):

substitute 2 instead of x here : 4x - 3

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes and now you need to find g(5)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

is that right @PFEH.1999

OpenStudy (anonymous):

no, 10 - 1/3 = 30/3 - 1/3 = 29/3

OpenStudy (skullpatrol):